Atlas Mines has adopted a policy of increasing the annual dividend on its common stock at a constant rate of 2.75 percent annually. The firm just paid an annual dividend of $1.67. What will the dividend be six years from now?
Answer #1

D6 = D0(1+G)^6

D6 = 1.67(1+0.0275)^6

D6 = 1.67*1.1767

D6 = 1.9652

D0           1.67
D1           1.72
D2           1.76
D3           1.81
D4           1.86
D5           1.91
D6           1.97